OpenStudy (anonymous):

Can someone help me with simplifying logs or exponential functions? the first one is: 2ln Radical(xe^x)

OpenStudy (anonymous):

\[2\ln(\sqrt{xe^x})\]

OpenStudy (anonymous):

the two out front comes up as an exponent so you get \[\ln(xe^x)=\ln(x)+\ln(e^x)=\ln(x)+x\]

OpenStudy (anonymous):

oh okayyy. can you help me with another possibly?

OpenStudy (anonymous):

go ahead and post! gotta run in three minutes